# Fine-Tuning Masked Diffusion for Provable Self-Correction

### v3

@@ Abstract
A natural desideratum for generative models is self-correction–detecting and revising low-quality tokens at inference. While Masked Diffusion Models (MDMs) have emerged as a promising approach for generative modeling in discrete spaces, their capacity for self-correction remains poorly understood. Prior attempts to incorporate self-correction into MDMs either require overhauling MDM architectures/training or rely on imprecise proxies for token quality, limiting their applicability. Motivated by this, we introduce PRISM–Plug-in Remasking for Inference-time Self-correction of Masked Diffusions–a lightweight, model-agnostic approach that applies to any pretrained MDM. Theoretically, PRISM defines a self-correction loss that provably learns per-token quality scores, without RL or a verifier. These quality scores are computed in the same forward pass with MDM and used to detect low-quality tokens. Empirically, PRISM advances MDM inference across domains and scales: Sudoku; unconditional text (170M); and code with LLaDA (8B).
@@ 1 Introduction
Masked Diffusion Models (MDMs) (sahoo2024simple; gat2024discrete; shi2024simplified) have emerged as a promising approach for generative modeling in discrete domains. At inference time, they start from a fully masked sequence and gradually unmask the masked tokens in arbitrary order to obtain a clean sequence. This flexibility in generation order, combined with recent approaches to deployment at scale (nie2025large; dream2025; gemini2025diffusion; labs2025mercury; gong2025diffucoder), has enabled them to surpass their autoregressive counterparts on several downstream tasks, such as reasoning, coding, and planning.
However, the standard MDM design lacks the ability to self-correct–a natural desideratum for generative models–identifying low-quality or incorrect tokens and correcting them at inference. In MDMs, once a mask token is unmasked, it remains fixed for the entire inference (Figure 1, left). Consequently, the model cannot revise early mistakes, a limitation several recent methods have tried to overcome (campbell2022continuous; lezama2022discrete; zhao2024informed; wang2025remasking; von2025generalized; peng2025pathplanningmaskeddiffusion). Unfortunately, existing approaches come with one of two drawbacks: they either (1) indirectly or inaccurately learn a notion of token quality that lacks theoretical grounding and is potentially inefficient to evaluate or (2) entirely change the MDM framework or architecture, limiting their applicability.
@@
Theoretically, PRISM introduces a novel self-correction loss that provably learns the per-token quality scores. It attaches a lightweight adapter to any pretrained MDM and fine-tunes with this loss; at inference, the learned scores detect low-quality tokens to remask and potentially be revised in subsequent inference steps (Figure 1, right). Practically, PRISM improves MDM inference across a range of domains and scales: Sudoku, unconditional text generation with a 170M MDM, and code generation with LLaDA (nie2025large), an open-source 8B MDM. Notably, on LLaDA, with <500<500 GPU-hours of fine-tuning, it acquires the ability to self-correct, outperforming baselines on MBPP austin2021program.
Why self-correction? Although self-correction is a common challenge in generative models where a token cannot be revised, it plays a more critical role in MDMs. This is because of a unique parallel sampling nature of MDMs, which significantly reduces inference cost. However, this parallel generation with only a few inference steps makes them more susceptible to dependency errors across positions (liu2024discrete; xu2024energy; kang2025parallelbench), as MDMs model per-position unmasking posteriors rather than the full joint distribution over positions. This limitation naturally motivates the need for self-correction mechanisms that can detect and rectify erroneous predictions. Indeed, our experiments show that PRISM becomes increasingly effective as more tokens are sampled in parallel, i.e., when MDM inference is performed with fewer sampling steps.
@@
Figure 1: PRISM overview. MDMs learn an unmasking posterior to unmask tokens, after which they remain fixed (Left). PRISM fine-tuning introduces per-token quality, which is used to detect incorrect tokens and remask them. At inference, the fine-tuned MDM jointly computes unmasking posterior and per-token quality, respectively performing unmasking and remasking (Right).
@@ 2.1 Masked Diffusion Models
Notation. Assume we aim to learn the distribution over sequences of length $L$ with vocabulary $\mathcal{V}$, e.g., natural language, namely the true distribution ${\mathbf{x}}\sim p_{\mathrm{data}}$. ${\mathbf{x}}^{i}$ denotes the $i$-th element of a given sequence ${\mathbf{x}}=({\mathbf{x}}^{1},\dots,{\mathbf{x}}^{L})$ and $\Delta(\mathcal{V})$ indicates the simplex of probability distributions over $\mathcal{V}$.
MDM training. Although MDMs can be interpreted from several viewpoints, we adopt an any-order language model viewpoint (ou2024your; zheng2024masked), which is essential to understanding our PRISM framework. Roughly speaking, MDM learns the posterior marginals over clean tokens in each masked position, conditioned on a partially masked sequence. We now formalize how this objective is defined and trained. To learn the posterior, MDM performs a masking process using an auxiliary mask token $\mathbf{m}$. For a given ${\mathbf{x}}\sim p_{\mathrm{data}}$, this masking process samples a partially masked sequence ${\mathbf{z}}\in(\mathcal{V}\cup\{\mathbf{m}\})^{L}$ defined in two equivalent ways (see Appendix B.2 for equivalence):
@@
which is the cross-entropy loss summed over all masked indices. Here, the expectation is taken over the joint distribution of $({\mathbf{x}},{\mathbf{z}})$ defined above. As desired, the unique minimizer $f_{\theta^{\star}}$ of the loss $\mathcal{L}_{\theta}$ satisfies $f_{\theta^{\star}}^{i}(v\,|\,{\mathbf{z}})=p({\mathbf{x}}^{i}=v\,|\,{\mathbf{z}})$. This connects the MDM training loss to any-order language models such as BERT (devlin2019bert), which model the posterior marginals at all masked positions. We note that the training loss equation 1 allows other equivalent formulations, as developed in previous MDM works (nie2025large; sahoo2024simple; shi2024simplified).
@@
Notably, step (a)—the choice of $\mathcal{S}$—is highly flexible and central to MDM’s gains on downstream tasks; $\mathcal{S}$ may be chosen arbitrarily. Two predominant strategies are independently including each masked index ${\mathbf{x}}_{t_{\ell}}^{i}=\mathbf{m}$ with a fixed probability (sahoo2024simple; shi2024simplified) or adaptively selecting indices, e.g., the most confident tokens or the leftmost tokens (chang2022maskgit; zheng2023reparameterized; kim2025train; nie2025large).
@@ 2.2 Self-correction for MDM via per-token quality
An ideal generative model would detect low-quality tokens at inference time and correct them when appropriate. Prior work has explored self-correction in LLMs, including kumar2024training; gou2023critic, and, recently, in MDMs. Despite varying instantiations, these methods for MDM can be organized into a common two-step procedure. At each inference step, given ${\mathbf{x}}_{t_{\ell}}$:
@@
Challenge. The key challenges are twofold: formulating a principled definition of per-token quality and constructing a model that estimates it. These can be formalized as follows: A per-token quality is a function $g_{\star}$ that takes any partially masked sequence $\mathbf{y}$} and returns position-wise scalars for non-masked positions ${\mathbf{y}}^{i}\neq\mathbf{m}$:
@@
  • Training-free approaches: DFM (gat2024discrete), ReMDM (wang2025remasking), P2-Self (peng2025pathplanningmaskeddiffusion) are drop-in and can be applied to any pretrained MDM, however they lack a precise target: scores are random (DFM, ReMDM), evaluated on a different input than ${\mathbf{x}}_{t_{\ell}}$ (ReMDM-conf), or inferred from logits that were not trained to present the token quality (P2-Self).
  • Methods that require rehauling MDM design: GIDD (von2025generalized), Informed Corrector (zhao2024informed), indeed articulate a valid target for self-correction but achieve it by changing the pretraining design or architecture, requiring full retraining.
  • Proxy-$\hat{{\mathbf{x}}}$ approaches: Token Critic (gou2023critic) and P2-Train (peng2025pathplanningmaskeddiffusion) train an external scorer to approximate the marginal likelihoods of a sampled clean sequence $\hat{{\mathbf{x}}}$. For a given ${\mathbf{x}}_{t_{\ell}}$, it first samples a reconstructed clean sequence $\hat{{\mathbf{x}}}$ via $f_{\theta}$ in a single step, then calculates the quality score via the trained module. Since the pretrained MDM has learned the per-token unmasking posterior, not the joint posterior across positions, $\hat{{\mathbf{x}}}$ can deviate a lot from a valid sample from the posterior. This method also requires an extra forward pass for scoring.
@@ 3 PRISM
In this section, we introduce Plug-in Remasking for Inference-time Self-correction of Masked diffusion models (PRISM). We substantiate its theoretical foundation in Section 3.1, present our empirical modeling design in Section 3.2, and explain how PRISM is used during inference in Section 3.3. Recall the definition of per-token quality from Section 2.2:
$g_{\star}\colon(\mathcal{V}\cup\{\mathbf{m}\})^{L}\to[0,1]^{L},\quad{\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g^{i}_{\star}({\mathbf{y}})\approx(\text{Per-token quality})}=(\text{``how likely'' ${\mathbf{y}}^{i}$ is given ${\mathbf{y}}$}).$ (2)
@@
${\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\phi}^{i}({\mathbf{y}})}\approx p({\mathbf{x}}^{i}={\mathbf{y}}^{i}\,|\,{\mathbf{y}}\oplus\mathbf{m}_{i})\in[0,1],\quad{\color[rgb]{0.25390625,0.41015625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.25390625,0.41015625,0.8828125}f_{\theta}^{i}(\cdot\,|\,{\mathbf{z}})}\approx p({\mathbf{x}}^{i}=\cdot\,|\,{\mathbf{z}})\in\Delta(\mathcal{V}),$
@@
A naive workaround is a distillation-based approach, supervising $\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\phi}^{i}({\mathbf{y}})$} with the teacher signal $\color[rgb]{0.25390625,0.41015625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.25390625,0.41015625,0.8828125}f_{\theta}^{i}(\cdot\,|\,{\mathbf{y}}\oplus\mathbf{m}^{i})$}. This has two drawbacks: First, even in an infinite data and compute regime, we cannot achieve perfect $g_{\phi^{\star}}$ if the teacher model $f_{\theta}$ is imperfect. Second, it is data-inefficient: each training pair $({\mathbf{y}},i)$ requires a fresh evaluation of $f_{\theta}$ on ${\mathbf{y}}\oplus\mathbf{m}_{i}$. Consequently, one forward pass of $f$ supervises only a single index, i.e., covering $m$ clean token positions in a sequence requires $m$ passes.
To address this input mismatch, zhao2024informed leverage a hollow-transformer $g_{\phi^{\prime}}$, with the property $g_{\phi^{\prime}}^{i}({\mathbf{y}}\oplus\mathbf{m}_{i})\approx p({\mathbf{x}}^{i}={\mathbf{y}}^{i}\,|\,{\mathbf{y}}\oplus\mathbf{m}_{i})$, enabling the use of the same loss as MDM. However, this requires adopting a new architecture (with a tweak on the attention mechanism) that processes $\mathbf{y}$} while equivalently behaving as if position $i$ were masked, which entails retraining and limiting plug-and-play applicability to pretrained MDMs.

### v4

@@ Abstract
A natural desideratum for generative models is self-correction–detecting and revising low-quality tokens at inference. While Masked Diffusion Models (MDMs) have emerged as a promising approach for generative modeling in discrete spaces, their capacity for self-correction remains poorly understood. Prior attempts to incorporate self-correction into MDMs either require overhauling MDM architectures/training or rely on imprecise proxies for token quality, limiting their applicability. Motivated by this, we introduce PRISM–Plug-in Remasking for Inference-time Self-correction of Masked Diffusions–a lightweight, model-agnostic approach that applies to any pretrained MDM. Theoretically, PRISM defines a self-correction loss that provably learns per-token quality scores, without RL or a verifier. These quality scores are computed in the same forward pass with MDM and used to detect low-quality tokens. Empirically, PRISM advances MDM inference across domains and scales: Sudoku; unconditional text (170M); and code with LLaDA (8B). We open-source our codebase in (Link).
Machine Learning, ICML
@@ 1 Introduction
Masked Diffusion Models (MDMs) (Sahoo et al., 2024; Gat et al., 2024; Shi et al., 2024) have emerged as a promising approach for generative modeling in discrete domains. At inference time, they start from a fully masked sequence and gradually unmask the masked tokens in arbitrary order to obtain a clean sequence. This flexibility in generation order, combined with recent approaches to deployment at scale (Nie et al., 2025; Ye et al., 2025; DeepMind, 2025; Labs et al., 2025; Gong et al., 2025), has enabled them to surpass their autoregressive counterparts on several downstream tasks, such as reasoning, coding, and planning.
Figure 1: PRISM. (Left) MDM learns unmasking posterior to unmask tokens, after which they remain fixed. (Middle) PRISM fine-tuning introduces per-token quality, which is used to detect incorrect tokens and remask them. At inference, the fine-tuned MDM jointly computes unmasking posterior and per-token quality, respectively performing unmasking and remasking. (Right) PRISM training pair generation: sample a masked sequence $\mathbf{z}$} from $\mathbf{x}$}; from $\mathbf{z}$}, unmask a chosen position using the pretrained MDM to obtain $\mathbf{y}$}.
However, the standard MDM design lacks the ability to self-correct–a natural desideratum for generative models—that is, identifying low-quality or incorrect tokens and correcting them at inference. This limitation is particularly crucial in MDMs due to their parallel, few step inference nature. Parallelized inference with only a few steps makes MDMs more susceptible to dependency errors across positions (Liu et al., 2024a; Xu et al., 2024; Kang et al., 2025), as they model per-position unmasking posteriors rather than the full joint distribution over positions.
Several recent works have attempted to overcome this limitation in MDMs (Campbell et al., 2022; Lezama et al., 2022b; Zhao et al., 2024; Wang et al., 2025; von Rütte et al., 2025; Peng et al., 2025). However, these approaches come with one of two drawbacks: they either (1) indirectly or inaccurately learn a notion of token quality and are potentially inefficient to evaluate or (2) entirely change the MDM framework or architecture, limiting their applicability.
@@
Theoretically, PRISM introduces a novel self-correction loss that provably learns the per-token quality scores. It attaches a lightweight adapter to any pretrained MDM and fine-tunes with this loss; at inference, the learned scores detect low-quality tokens to remask and potentially be revised in subsequent inference steps (Figure 1, right). Practically, PRISM improves MDM inference across a range of domains and scales: Sudoku, unconditional text generation with a 170M MDM, and code generation with LLaDA (Nie et al., 2025), an open-source 8B MDM. Notably, on LLaDA, with <500<500 GPU-hours of fine-tuning, it acquires the ability to self-correct, outperforming baselines on MBPP (Austin et al., 2021b) and HumanEval (Chen, 2021).
@@
Concurrent work. Meshchaninov et al. (2025); Li et al. (2025); Schiff et al. (2026), while differing in technical details, also address the self-correction problem for MDMs by either training an additional module or fine-tuning the MDM backbone that takes $\hat{{\mathbf{x}}}$ as input, thus falling under the class of proxy-$\hat{{\mathbf{x}}}$ approaches described in Section 2.2. Most closely related, (Huang et al., 2025) shares the motivation and experimental setup. More recently, Bie et al. (2026) trains large-scale MDMs at 16B and 100B parameters. Although the details are not disclosed, their codebase suggests that they learn a per-position distribution rather than a scalar. We return to this discussion in Section 5.111This paragraph has been included since the second version of this manuscript.
Conflict of Interest Disclosure. The authors declare no financial conflicts of interest.
@@ 2.1 Masked Diffusion Models
Notation. Assume we aim to learn the distribution over sequences of length $L$ with vocabulary $\mathcal{V}$, namely the true distribution ${\mathbf{x}}\sim p_{\mathrm{data}}$. ${\mathbf{x}}^{i}$ denotes the $i$-th element of a given sequence ${\mathbf{x}}=({\mathbf{x}}^{1},\dots,{\mathbf{x}}^{L})$ and $\Delta(\mathcal{V})$ indicates the simplex of probability distributions over $\mathcal{V}$.
MDM training. Although MDMs can be interpreted from several viewpoints, we adopt an any-order language model viewpoint (Ou et al., 2024; Zheng et al., 2024), which is essential to understanding our PRISM framework. Roughly speaking, MDM learns the posterior marginals over clean tokens in each masked position, conditioned on a partially masked sequence. We now formalize how this objective is defined and trained. To learn the posterior, MDM performs a masking process using an auxiliary mask token $\mathbf{m}$. For a given ${\mathbf{x}}\sim p_{\mathrm{data}}$, this masking process samples a partially masked sequence ${\mathbf{z}}\in(\mathcal{V}\cup\{\mathbf{m}\})^{L}$ defined in two equivalent ways (see Appendix B.2 for equivalence):
@@
which is the cross-entropy loss summed over all masked indices. Here, the expectation is taken over the joint distribution of $({\mathbf{x}},{\mathbf{z}})$ defined above. As desired, the unique minimizer $f_{\theta^{\star}}$ of the loss $\mathcal{L}_{\theta}$ satisfies $f_{\theta^{\star}}^{i}(v\,|\,{\mathbf{z}})=p({\mathbf{x}}^{i}=v\,|\,{\mathbf{z}})$. This connects the MDM training loss to any-order language models such as BERT (Devlin et al., 2019), which model the posterior marginals at all masked positions. We note that the training loss (1) allows other equivalent formulations, as developed in previous MDM works (Nie et al., 2025; Sahoo et al., 2024; Shi et al., 2024).
@@
Notably, step (a)—the choice of $\mathcal{S}$—is highly flexible and central to MDM’s gains on downstream tasks; $\mathcal{S}$ may be chosen arbitrarily. Two predominant strategies are independently including each masked index ${\mathbf{x}}_{t_{\ell}}^{i}=\mathbf{m}$ with a fixed probability (Sahoo et al., 2024; Shi et al., 2024) or adaptively selecting indices, e.g., the most confident tokens or the leftmost tokens (Chang et al., 2022; Zheng et al., 2023; Kim et al., 2025; Nie et al., 2025).
@@ 2.2 Self-correction for MDM via Per-token Quality
An ideal generative model would detect low-quality tokens at inference time and correct them when appropriate. Prior work has explored self-correction in LLMs, including (Kumar et al., 2024; Gou et al., 2023), and, recently, in MDMs. Despite varying instantiations, these methods for MDM can be organized into a two-step recipe. At each inference step, given ${\mathbf{x}}_{t_{\ell}}$:
@@
Challenge. The key challenges are twofold: formulating a principled definition of per-token quality and constructing a model that estimates it. These can be formalized as follows: A per-token quality is a function $g_{\star}$ that takes any partially masked sequence $\mathbf{y}$} and returns position-wise scalars for non-masked positions ${\mathbf{y}}^{i}\neq\mathbf{m}$: $g_{\star}\colon(\mathcal{V}\cup\{\mathbf{m}\})^{L}\to[0,1]^{L}$,
${\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g^{i}_{\star}({\mathbf{y}})\approx(\text{Per-token quality})}=(\text{\emph{how likely} ${\mathbf{y}}^{i}$ is given ${\mathbf{y}}$}).$
@@
  • Training-free approaches: DFM (Gat et al., 2024), ReMDM (Wang et al., 2025), P2-Self (Peng et al., 2025) are drop-in and can be applied to any pretrained MDM, however they lack a precise target: scores are random (DFM, ReMDM), evaluated on a different input than ${\mathbf{x}}_{t_{\ell}}$ (ReMDM-conf), or inferred from logits that were not trained to present the token quality (P2-Self).
  • Methods that require rehauling MDM design: GIDD (von Rütte et al., 2025), Informed Corrector (Zhao et al., 2024), indeed articulate a valid target for self-correction but achieve it by changing the pretraining design or architecture, requiring full retraining.
  • Proxy-$\hat{{\mathbf{x}}}$ approaches: Token Critic (Gou et al., 2023) and P2-Train (Peng et al., 2025) train an external scorer to approximate the marginal likelihoods of a sampled clean sequence $\hat{{\mathbf{x}}}$. For a given ${\mathbf{x}}_{t_{\ell}}$, it first samples a reconstructed clean sequence $\hat{{\mathbf{x}}}$ via $f_{\theta}$ in a single step, then calculates the quality score via the trained module. Since the pretrained MDM has learned the per-token unmasking posterior, not the joint posterior across positions, $\hat{{\mathbf{x}}}$ can deviate a lot from a valid sample from the posterior. This method also requires an extra forward pass for scoring.
@@ 3 PRISM
In this section, we introduce Plug-in Remasking for Inference-time Self-correction of Masked diffusion models (PRISM). We substantiate its theoretical foundation in Section 3.1, present our empirical modeling design in Section 3.2, and explain how PRISM is used during inference in Section 3.3. Recall the definition of per-token quality from Section 2.2: $g_{\star}\colon(\mathcal{V}\cup\{\mathbf{m}\})^{L}\to[0,1]^{L}$,
${\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g^{i}_{\star}({\mathbf{y}})\approx(\text{Per-token quality})}=(\text{\emph{how likely} ${\mathbf{y}}^{i}$ is given ${\mathbf{y}}$}).$
@@
Algorithm 1 PRISM fine-tuning from a pretrained MDM
1:Require: Pretrained MDM backbone $f_{\theta}$, per-token quality head $g_{\theta}$, true distribution $p_{\mathrm{data}}$, regularization constant $\lambda>0$, number of unmasking tokens $k\in\mathbb{N}$.
3:# Sample $\mathbf{y}$} from $\mathbf{z}$} with $f_{\mathrm{sg}(\theta)}$
4: Select a subset $\mathcal{S}\subseteq\{i\mid{\mathbf{z}}^{i}=\mathbf{m}\}$ with $|\mathcal{S}|=k$.
5: Obtain $\mathbf{y}$} by replacing ${\mathbf{z}}^{i}=\mathbf{m}$ to $v^{i}\sim{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}f_{\mathrm{sg}(\theta)}^{i}(\cdot\,|\,{\mathbf{z}})}$ for each $i\in\mathcal{S}$.
6:# Calculate Loss
[…]
@@
${\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\phi}^{i}({\mathbf{y}})}\approx p({\mathbf{x}}^{i}={\mathbf{y}}^{i}\,|\,{\mathbf{y}}\oplus\mathbf{m}_{i})\in[0,1],$
${\color[rgb]{0.25390625,0.41015625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.25390625,0.41015625,0.8828125}f_{\theta}^{i}(\cdot\,|\,{\mathbf{z}})}\approx p({\mathbf{x}}^{i}=\cdot\,|\,{\mathbf{z}})\in\Delta(\mathcal{V}),$
@@
A naive workaround is a distillation-based approach, supervising $\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\phi}^{i}({\mathbf{y}})$} with the teacher signal $\color[rgb]{0.25390625,0.41015625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.25390625,0.41015625,0.8828125}f_{\theta}^{i}(\cdot\,|\,{\mathbf{y}}\oplus\mathbf{m}_{i})$}. This has two drawbacks: First, even in an infinite data and compute regime, we cannot achieve perfect $g_{\phi^{\star}}$ if the teacher model $f_{\theta}$ is imperfect. Second, it is data-inefficient: each training pair $({\mathbf{y}},i)$ requires a fresh evaluation of $f_{\theta}$ on ${\mathbf{y}}\oplus\mathbf{m}_{i}$. Consequently, one forward pass of $f$ supervises only a single index, i.e., covering $m$ clean token positions in a sequence requires $m$ passes.
To address this input mismatch, Zhao et al. (2024) leverage a hollow-transformer $g_{\phi^{\prime}}$, with the property $g_{\phi^{\prime}}^{i}({\mathbf{y}}\oplus\mathbf{m}_{i})\approx p({\mathbf{x}}^{i}={\mathbf{y}}^{i}\,|\,{\mathbf{y}}\oplus\mathbf{m}_{i})$, enabling the use of the same loss as MDM. However, this requires adopting a new architecture (with a tweak on the attention mechanism) that processes $\mathbf{y}$} while equivalently behaving as if position $i$ were masked, which entails retraining and limiting plug-and-play applicability to pretrained MDMs.